OpenStudy (anonymous):

how do i simplify (3a^5)^-2/a^4

OpenStudy (asapbleh):

Since you are trying to simplify the fraction that contains an negative exponent, its rules are this: when x^-2 = 1/x^2 when (3+x)^-2 = 1/(3+x)^2 because the whole thing inside the parentheses is squared, so you much move the whole thing down.

OpenStudy (asapbleh):